OpenStudy (anonymous):

write -2x^2(-5x^2+4x^3) in standard form ?? explain?? :)

OpenStudy (agent0smith):

First you need to distribute/simplify, then worry about the standard form

OpenStudy (anonymous):

i haven't done this in awhile but i know i distribute the -2x^2 to the paranthesis but do i add -2 to -5 so that becomes -7x^4 but idk

OpenStudy (anonymous):

so lost :S

OpenStudy (agent0smith):

You're distributing - this means you need to multiply everything a(b+c) = ab + ac eg 2x(4x+3) = \[\Large 2x(4x+3) = 2*4*x*x + 2*3*x = 8x^2 + 6x\]

OpenStudy (anonymous):

thats not any of the answers to the question D:

OpenStudy (agent0smith):

That was an example :/

OpenStudy (anonymous):

OH omfg i feel so dumb

OpenStudy (anonymous):

im working multiple problems rn but hold on ill try and distribute

OpenStudy (anonymous):

it is 10x^4 + 8x^5 ??

OpenStudy (anonymous):

or -8x^5

OpenStudy (agent0smith):

^correct, the negative

OpenStudy (agent0smith):

Now standard form just means you need to change the order, put the highest power of x first.